# Beyond QA Pairs: Assessing Parameter-Efficient Fine-Tuning for Fact Embedding in LLMs

### v1

@@ Appendix A Appendix
Evaluator LLM Prompt Task Description: An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given. 1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general. 2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric. 3. The output format should look as follows: Feedback: {{write a feedback for criteria}} [RESULT] {{an integer number between 1 and 5}} 4. Please do not generate any other opening, closing, and explanations. Be sure to include [RESULT] in your output. The instruction to evaluate: {instruction} Response to evaluate: {response} Reference Answer (Score 5): {reference_answer} Score Rubrics: Is the response correct, accurate, and factual based on the reference answer? Score 1: The response is completely incorrect, inaccurate, and/or not factual. Score 2: The response is mostly incorrect, inaccurate, and/or not factual. Score 3: The response is somewhat correct, accurate, and/or factual. Score 4: The response is mostly correct, accurate, and/or factual. Score 5: The response is completely correct, accurate, and factual. Feedback:
@@
D-RAG and D-Naive Example Question: How does <COMPANY> Fabric simplify the configuration of networking for a Metal-hosted cluster? D-RAG Answer: <COMPANY> Fabric simplifies the configuration of networking for a Metal-hosted cluster by providing a global software-defined interconnection platform. It allows you to connect the Metal site to remote cloud or other infrastructure easily. With <COMPANY> Fabric, you can set up a fabric-based networking configuration, which not only makes the setup process easy but also allows you to keep your inter-cluster traffic private if desired. Additionally, <COMPANY> Fabric supports a container network interface (CNI) plugin designed for multi-cluster networking, making it relatively easy to configure. This simplifies the networking setup for Metal-hosted clusters and provides flexibility for future changes to the Kubernetes environment or architecture. D-Naive Answer: <COMPANY> Fabric simplifies the configuration of networking for a Metal-hosted cluster by providing a global software-defined interconnection platform. By using <COMPANY> Fabric, you can easily connect the Metal site to other remote cloud or infrastructure where your other cluster is hosted. This allows for seamless communication between clusters and eliminates the complexities of configuring networking tools that may vary from site to site. Additionally, <COMPANY> Fabric enables you to keep your inter-cluster traffic private if desired.
@@
Call Transcript and Next Step Example Sales: Hi, John here from the sales team. I wanted to discuss a new feature idea for our product. Developer: Hi John, sure. What’s the feature you have in mind? Sales: I was thinking of adding a chatbot feature that can assist customers with their inquiries. Developer: That sounds interesting. How do you envision the chatbot to work? Sales: The chatbot should be able to answer frequently asked questions, provide product information, and guide customers through the purchasing process. Developer: Okay. Are there any specific platforms or technologies you would like us to use for building the chatbot? Sales: We’re open to suggestions, but it would be great if the chatbot can be integrated into our existing website and mobile app. Developer: Got it. I’ll look into the available chatbot frameworks and APIs that we can leverage for this project. Sales: Perfect. Let’s also consider personalization options and the ability to hand over to a human agent if needed. Developer: Definitely, we can add those features as well. I’ll start working on the requirements and come up with a development plan. Sales: Great! Once we have the chatbot ready, we can demonstrate it to potential customers and gather feedback. Developer: That’s a good plan. We’ll need to involve the marketing team to create awareness and promote the new feature. Sales: Absolutely. I’ll reach out to the marketing team and coordinate with them to align our strategies. Next Steps: Sales: Coordinate with the marketing team to align strategies and create awareness about the new chatbot feature. Developer: Research chatbot frameworks and APIs, and come up with a development plan for the chatbot feature.
@@
Product description and Pitch Product Description: <COMPANY> Metal is a global platform that allows users to easily provision bare metal servers anywhere in the world. It provides state-of-the-art hardware without any virtualization layer, offering pure, unadulterated iron. With <COMPANY> Metal, users can experience the ease and convenience of deploying servers from the comfort of their laptop. Product Pitch: Experience the power of <COMPANY> Metal, the global platform for bare metal server provisioning. With state-of-the-art hardware and easy deployment process, <COMPANY> Metal allows you to spin up servers anywhere in the world. Say goodbye to virtualization layers and enjoy the pure performance of unadulterated iron.
@@
Product Recommendation For the following customer requirement and given list of product description, output the name of the product which can be recommended to customer to solve their problem Customer requirement: The customer requires a networking solution that enables seamless communication between clusters in a multi-cluster Kubernetes environment, running at different sites. Products: <COMPANY> Metal offers bare metal servers that are ready to use when and where needed. They provide the flexibility and reliability required to navigate challenging supply chains and ensure on-time delivery of IT hardware. Network Edge is a service offered by <COMPANY> that allows users to create virtual devices with primary and secondary redundancy. It provides a flexible and scalable networking solution for businesses. The Unified Cross Connects Portal is a platform provided by <COMPANY> that allows customers to manage and order Cross Connects with ease. It provides a centralized interface for accessing Cross Connect services, enabling customers to efficiently configure and schedule their connections. <COMPANY> Fabric is a global software-defined interconnection platform provided by <COMPANY>. It allows for easy networking configuration in a multi-cluster Kubernetes environment, enabling seamless communication between clusters running at different sites. With <COMPANY> Fabric, you can ensure secure and private inter-cluster traffic while simplifying your networking setup. <COMPANY> Metal is a cloud infrastructure offering that allows users to manage Metal resources in event-driven configurations. It provides a growing Ansible collection as a provider for seamless integration. Output: <COMPANY> Fabric

### v2

@@ Appendix A Appendix
Evaluator LLM Prompt Task Description: {internallinenumbers*} An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given. {internallinenumbers*} 1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general. {internallinenumbers*} 2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric. {internallinenumbers*} 3. The output format should look as follows: Feedback: {{write a feedback for criteria}} [RESULT] {{an integer number between 1 and 5}} {internallinenumbers*} 4. Please do not generate any other opening, closing, and explanations. Be sure to include [RESULT] in your output. The instruction to evaluate: {instruction} Response to evaluate: {response} Reference Answer (Score 5): {reference_answer} Score Rubrics: {internallinenumbers*} Is the response correct, accurate, and factual based on the reference answer? {internallinenumbers*} Score 1: The response is completely incorrect, inaccurate, and/or not factual. Score 2: The response is mostly incorrect, inaccurate, and/or not factual. Score 3: The response is somewhat correct, accurate, and/or factual. Score 4: The response is mostly correct, accurate, and/or factual. Score 5: The response is completely correct, accurate, and factual. Feedback:
@@
D-RAG and D-Naive Example {internallinenumbers*} Question: How does <COMPANY> Fabric simplify the configuration of networking for a Metal-hosted cluster? {internallinenumbers*} D-RAG Answer: <COMPANY> Fabric simplifies the configuration of networking for a Metal-hosted cluster by providing a global software-defined interconnection platform. It allows you to connect the Metal site to remote cloud or other infrastructure easily. With <COMPANY> Fabric, you can set up a fabric-based networking configuration, which not only makes the setup process easy but also allows you to keep your inter-cluster traffic private if desired. Additionally, <COMPANY> Fabric supports a container network interface (CNI) plugin designed for multi-cluster networking, making it relatively easy to configure. This simplifies the networking setup for Metal-hosted clusters and provides flexibility for future changes to the Kubernetes environment or architecture. {internallinenumbers*} D-Naive Answer: <COMPANY> Fabric simplifies the configuration of networking for a Metal-hosted cluster by providing a global software-defined interconnection platform. By using <COMPANY> Fabric, you can easily connect the Metal site to other remote cloud or infrastructure where your other cluster is hosted. This allows for seamless communication between clusters and eliminates the complexities of configuring networking tools that may vary from site to site. Additionally, <COMPANY> Fabric enables you to keep your inter-cluster traffic private if desired.
@@
Call Transcript and Next Step Example {internallinenumbers*} Sales: Hi, John here from the sales team. I wanted to discuss a new feature idea for our product. Developer: Hi John, sure. What’s the feature you have in mind? {internallinenumbers*} Sales: I was thinking of adding a chatbot feature that can assist customers with their inquiries. Developer: That sounds interesting. How do you envision the chatbot to work? {internallinenumbers*} Sales: The chatbot should be able to answer frequently asked questions, provide product information, and guide customers through the purchasing process. {internallinenumbers*} Developer: Okay. Are there any specific platforms or technologies you would like us to use for building the chatbot? {internallinenumbers*} Sales: We’re open to suggestions, but it would be great if the chatbot can be integrated into our existing website and mobile app. {internallinenumbers*} Developer: Got it. I’$l$ look into the available chatbot frameworks and APIs that we can leverage for this project. {internallinenumbers*} Sales: Perfect. Let’s also consider personalization options and the ability to hand over to a human agent if needed. {internallinenumbers*} Developer: Definitely, we can add those features as well. I’$l$ start working on the requirements and come up with a development plan. {internallinenumbers*} Sales: Great! Once we have the chatbot ready, we can demonstrate it to potential customers and gather feedback. {internallinenumbers*} Developer: That’s a good plan. We’$l$ need to involve the marketing team to create awareness and promote the new feature. {internallinenumbers*} Sales: Absolutely. I’$l$ reach out to the marketing team and coordinate with them to align our strategies. Next Steps: {internallinenumbers*} Sales: Coordinate with the marketing team to align strategies and create awareness about the new chatbot feature. {internallinenumbers*} Developer: Research chatbot frameworks and APIs, and come up with a development plan for the chatbot feature.
@@
Product description and Pitch Product Description: {internallinenumbers*} <COMPANY> Metal is a global platform that allows users to easily provision bare metal servers anywhere in the world. It provides state-of-the-art hardware without any virtualization layer, offering pure, unadulterated iron. With <COMPANY> Metal, users can experience the ease and convenience of deploying servers from the comfort of their laptop. Product Pitch: {internallinenumbers*} Experience the power of <COMPANY> Metal, the global platform for bare metal server provisioning. With state-of-the-art hardware and easy deployment process, <COMPANY> Metal allows you to spin up servers anywhere in the world. Say goodbye to virtualization layers and enjoy the pure performance of unadulterated iron.
@@
Product Recommendation {internallinenumbers*} For the following customer requirement and given list of product description, output the name of the product which can be recommended to customer to solve their problem Customer requirement: {internallinenumbers*} The customer requires a networking solution that enables seamless communication between clusters in a multi-cluster Kubernetes environment, running at different sites. Products: {internallinenumbers*} <COMPANY> Metal offers bare metal servers that are ready to use when and where needed. They provide the flexibility and reliability required to navigate challenging supply chains and ensure on-time delivery of IT hardware. {internallinenumbers*} Network Edge is a service offered by <COMPANY> that allows users to create virtual devices with primary and secondary redundancy. It provides a flexible and scalable networking solution for businesses. {internallinenumbers*} The Unified Cross Connects Portal is a platform provided by <COMPANY> that allows customers to manage and order Cross Connects with ease. It provides a centralized interface for accessing Cross Connect services, enabling customers to efficiently configure and schedule their connections. {internallinenumbers*} <COMPANY> Fabric is a global software-defined interconnection platform provided by <COMPANY>. It allows for easy networking configuration in a multi-cluster Kubernetes environment, enabling seamless communication between clusters running at different sites. With <COMPANY> Fabric, you can ensure secure and private inter-cluster traffic while simplifying your networking setup. {internallinenumbers*} <COMPANY> Metal is a cloud infrastructure offering that allows users to manage Metal resources in event-driven configurations. It provides a growing Ansible collection as a provider for seamless integration. Output: <COMPANY> Fabric